OpenStudy (anonymous):

Line AEB and line CED intersects at E. If measure of angle Bec=70, find the measure of angle AED, measure of angle DEB, and the measure of angle AEC

OpenStudy (anonymous):

do you have a pic?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Angle AED = Angle BED = 70 (vertically opposite angles) Angle DEB = 180 - (Angle BEC) = 180-70 = 110 (supplementary angles ie. sum = 180) Angle AEC = Angle DEB = 110 (vertically opposite angles)